# Efficient Compression of Redshift-Space Distortion Data for Late-Time Modified Gravity Models

### v1

@@ I Introduction
The accelerated expansion of the universe is a very well-established observational fact, supported by a wide range of diverse cosmological measurements Aghanim et al. (2020); Alam et al. (2021); Abbott et al. (2022); Brout et al. (2022). While in the context of the standard model of cosmology, a.k.a $\Lambda$CDM model, this acceleration is explained by means of a cosmological constant $\Lambda>0$ with associated positive energy density $\rho_{\Lambda}=\Lambda/8\pi G$ and negative pressure $p_{\Lambda}=-\rho_{\Lambda}$ (Peebles and Ratra, 2003; Padmanabhan, 2003), cosmological data still leaves room for dynamical forms of dark energy (DE) different from a rigid (immutable and homogeneous) $\Lambda$ Amendola and Tsujikawa (2015); Solà Peracaula (2022). In fact, from a phenomenological perspective, deviations from the standard model might be necessary in view of the tensions afflicting it, such as the Hubble and growth tensions (see, e.g., the reviews Perivolaropoulos and Skara (2022); Abdalla et al. (2022) and references therein). They seem to require modifications to the background dynamics and/or the perturbation equations of $\Lambda$CDM.
@@
Data on large-scale structure (LSS) have played an important role in constraining cosmological models and may be relevant for consolidating the aforementioned hints of dynamical DE. The data on redshift-space distortions (RSD) and weak gravitational lensing tend to favor a lower growth of matter perturbations in the late-time universe compared to the Planck best-fit $\Lambda$CDM model Macaulay et al. (2013); Joudaki et al. (2018); Gómez-Valent and Solà (2017); Nesseris et al. (2017); Gómez-Valent and Solà Peracaula (2018); Wright et al. (2020); Nunes and Vagnozzi (2021); Nguyen et al. (2023); Adil et al. (2023). Interestingly, upcoming RSD data from DESI will cover a redshift range up to $z\sim 2.1$, extending by $\sim 0.6$ units the range covered by past surveys, and with small relative uncertainties of $3-6\%$ in the intermediate range $z\in(0.4,1.6)$ Adame et al. (2024b). These new data will allow us to better assess the status of the growth tension.
@@
In this work, we aim to compress the information contained in RSD data into variables that parameterize deviations from the standard growth of perturbations in the late universe, with the aid of a couple of sensible priors. These variables should have a crystal-clear physical interpretation and be applicable to a broad range of modified gravity and dark energy models. More concretely, in this paper, we model the gravitational coupling as a step function and introduce 2 or 3 parameters (depending on the case under study) representing the step widths of the functions. We demonstrate that this model can reproduce the growth of density fluctuations of effective field theories of DE (EFTofDE) with observational accuracy. Other models, such as Galileon gravity, Dvali-Gabadadze-Porrati (DGP) gravity, and $f(R)$ gravity, have already been examined in Denissenya and Linder (2017). We apply the compression technique proposed in Denissenya and Linder (2017) and study its robustness and universality by applying it to three different models within the large class of EFTofDE. In most works in the literature, the time dependence of the EFT parameters or the effective gravitational coupling is assumed when deriving or forecasting constraints from observations Zhao et al. (2009); Ade et al. (2016); Ferté et al. (2019); Abbott et al. (2019); Sakr and Martinelli (2022); Abbott et al. (2023); Casas et al. (2023). These constraints strongly depend on the assumed functional form. More general (continuous) reconstructions of the gravitational coupling can be performed using more sophisticated techniques, such as a Bayesian reconstruction or Gaussian Processes Raveri et al. (2023); Pogosian et al. (2022); Ruiz-Zapatero et al. (2022); Mu et al. (2023), but the resulting constraints cannot be easily projected onto constraints on the parameters of specific models. See also Zheng et al. (2024) for a very model-independent method to constrain deviations from standard growth that does not make use of RSD data.
@@
The 2-3 parameter description proposed in Denissenya and Linder (2017) offers a quite model-independent way to reveal the physical characteristics of the underlying theory required to explain the RSD data. It relies on the cosmological principle and the covariant conservation of matter. It can account, e.g., for phenomena such as the suppression of growth at late times Macaulay et al. (2013); Joudaki et al. (2018); Gómez-Valent and Solà (2017); Nesseris et al. (2017); Gómez-Valent and Solà Peracaula (2018); Wright et al. (2020); Nunes and Vagnozzi (2021); Nguyen et al. (2023); Adil et al. (2023). Additionally, we will see that it provides an efficient method for constraining EFT parameters with specific time dependencies by projecting the compressed information onto the latter.
@@ IV.2 Current LSS data
In Table 1, we present the list with the 15 RSD data points employed in Sec. VI.1, together with the corresponding references. They are provided by different galaxy surveys and cover the redshift range $z\in(0.01,1.5)$.
@@ V.1 Methodology
    The latter is obtained in $\Lambda$CDM using the DESI BAO data, the CMB temperature and polarization data from Planck and the CMB lensing data from Planck and ACT Adame et al. (2024a). The results of the fitting analyses carried out using Eq. (15) are displayed in Tables 3-6.
@@
    with $A_{i},B_{i},D_{i}$ fitting parameters, is accurate enough. We illustrate this in Fig. 1. See also Sec. V.2.
@@
    where $\boldsymbol{p}(c_{j},\Omega_{m})$ and $\boldsymbol{p}_{\mathrm{mean}}$ are the vectors containing the various $\mu_{i}$’s and $\Omega_{m}$ from theory and data, respectively. The covariance matrix $Cov$ is also obtained in point 2. With a prior on $\Omega_{m}$, we found that the Gaussian approximation for these parameters is valid for the DESI mock data.
@@ V.2 Results
Figure 1: Comparison of the theoretical curves of $f(z)\sigma_{8}(z)$ obtained for the various EFT models employed in this paper with $c_{B}=0.4$ and $\Omega_{m}=0.3$ (in blue), and the corresponding best-fit curves of the bin model (in red or orange). The latter have been obtained following the procedure explained in point 3 of Sec. V.1. In the right column we plot the percentage relative difference $\Delta f\sigma_{8}(z)/f\sigma_{8}(z)$ [%] between the theoretical and best-fit results. They are smaller than $\sim 1.5\%$ at $z<0.5$ and smaller than $0.2\%$ at $z>0.5$ in absolute value, with the only exception of the fit of the propto-Scale model using $\mu$-2param. As explained in Sec. V.2.3, in this case we need 3 parameters to describe accurately enough the theory.
@@
This is the covariance matrix that enters Eq. (20) and is employed in the second Monte Carlo run to obtain the projected constraints on the EFT parameters in the bin method. While we present the mean values and errors of the $\mu_{i}$’s from all analyses in this work in the various tables, we provide the covariance matrix only for this particular case to maintain compactness.
@@ V.2.2 Inv-Hubble-Squared
with $j=B$ (if $c_{M}=0$) or $j=M$ (if $c_{B}=0$). We calculate the tension $T$ using our constraints at 68% C.L. We find that the tension $T<0.12$ in all cases, which again demonstrates the good performance of our method.
@@
Table 6: Same as Table 3, but for the two propto-Scale models studied in this paper. On the top we show the results obtained considering $c_{B}\neq 0$ and $c_{M}=0$, while on the bottom we display the results obtained when $c_{B}=-c_{M}$. We use a $\Lambda$CDM background.
@@ VI.1 Current data
$\mu_{I}(a)=1+\frac{1}{2}\left[\mu_{2}+\left(\frac{\mu_{1}-\mu_{2}}{2}\right)\left(1+\tanh\{\beta(2-1/a)\}\right)-1\right]\times\left[1+\tanh\{\beta(4-1/a)\}\right]\,.$ (23)
@@
Figure 3: Example of $\mu_{I}(z)$ with $\mu_{1}=1.3$ and $\mu_{2}=1.1$, and of $\mu_{II}(z)$ with $\mu_{1}=1.1$, $\mu_{2}=0.9$ and $\mu_{3}=1.2$. In both cases we have set $\beta=5$. See Eqs. (23) and (24).
$\mu_{II}(a)=1+\frac{1}{2}(\tilde{\mu}(a)-1)\left[1+\tanh\{\beta(10-1/a)\}\right]\,,$ (24)
@@
A couple of examples of $\mu_{I}(z)$ and $\mu_{II}(z)$ are provided in Fig. 3.
@@
Our results are provided in Table 7 and Fig. 4. In Fig. 5 we compare the best-fit curves of $f(z)\sigma_{8}(z)$ obtained with $\mu$-2param, $\mu$-3param and the $\Lambda$CDM. We find $\mu_{1}<1$ at $\sim 2.6-2.8\sigma$ C.L. with both parametrizations. The data on $f(z_{i})\sigma_{8}(z_{i})$ prefer a lower amount of structure formation in the universe than in $\Lambda$CDM at $z<1$. This explains why our parametrizations are capable of alleviating the tension and decreasing significantly the values of $\chi^{2}_{\rm min}$. Notice that 13 out of the 15 data points on $f(z_{i})\sigma_{8}(z_{i})$ employed in the analysis are inside that redshift range. These results are aligned with those reported in previous works in the literature, see e.g. Macaulay et al. (2013); Gómez-Valent and Solà (2017); Nesseris et al. (2017); Gómez-Valent and Solà Peracaula (2018); Nunes and Vagnozzi (2021); Nguyen et al. (2023) and references therein. At redshifts $z>1$ we only have two data points, with larger central values than in $\Lambda$CDM. This leads to values of $\mu_{2}$ in the parametrization $\mu$-2param slightly greater than $1$, $\mu_{2}=1.10\pm 0.09$, but compatible with the standard value ($\mu=1$) at $\sim 1\sigma$ C.L. In $\mu$-3param we find a significant degeneracy in the plane $\mu_{2}-\mu_{3}$ (cf. Fig. 4) due to the lack of data points at $z>1$. Current data are not able to set strong constraints on these two parameters simultaneously, and actually the posterior distributions hit the prior boundaries.
@@
Figure 5: Best-fit curves of $f(z)\sigma_{8}(z)$ obtained using the two parametrizations $\mu_{I}(z)$ and $\mu_{II}(z)$, Eqs. (23) and (24), together with the data points of Table 1 and the curve obtained using the $\Lambda$CDM best-fit model.
@@ VII Conclusions
Although we focused on obtaining the constraints on $\mu$ using the compressed observable $f\sigma_{8}$ in this paper, our method can also be eventually applied to data on $f\sigma_{12}$ Sánchez (2020); Semenaite et al. (2022, 2023), in case they become available in the future. The 2-3 parameter description of $\mu$ can be also implemented in the EFTofLSS approach to obtain the constraints on $\mu_{i}$ directly. It is known that the EFTofLSS approach suffers from prior projection effects due to a large number of nuisance parameters and it is advantageous to have a fewer parameters. The 2-3 parameter description is ideal as we do not need to assume a specific time dependence of $\mu$, and it is general enough to cover a wide variety of models as shown in this paper and Denissenya and Linder (2017). It remains to see whether the same binning of $\mu$ is enough for other LSS observables such as CMB (ISW, lensing) and weak lensing. This will be addressed in future work.

### v2

@@ I Introduction
The accelerated expansion of the universe is a very well-established observational fact, supported by a wide range of diverse cosmological measurements Aghanim et al. (2020); Alam et al. (2021); Abbott et al. (2022); Brout et al. (2022). While in the context of the standard model of cosmology, also known as Lambda Cold Dark Matter ($\Lambda$CDM) model, this acceleration is explained by means of a cosmological constant $\Lambda>0$ with associated positive energy density $\rho_{\Lambda}=\Lambda/8\pi G$ and negative pressure $p_{\Lambda}=-\rho_{\Lambda}$ (Peebles and Ratra, 2003; Padmanabhan, 2003), cosmological data still leaves room for dynamical forms of dark energy (DE) different from a rigid (immutable and homogeneous) $\Lambda$ Amendola and Tsujikawa (2015); Solà Peracaula (2022). In fact, from a phenomenological perspective, deviations from the standard model might be necessary in view of the tensions afflicting it, such as the Hubble and growth tensions (see, e.g., the reviews Perivolaropoulos and Skara (2022); Abdalla et al. (2022) and references therein). They seem to require modifications to the background dynamics and/or the perturbation equations of $\Lambda$CDM.
@@
Data on large-scale structure (LSS) have played an important role in constraining cosmological models and may be relevant for consolidating the aforementioned hints of dynamical DE. The data on redshift-space distortions (RSD) and weak gravitational lensing tend to favor a lower growth of matter perturbations in the late-time universe compared to the Planck best-fit $\Lambda$CDM model Macaulay et al. (2013); Joudaki et al. (2018); Gómez-Valent and Solà (2017); Nesseris et al. (2017); Gómez-Valent and Solà Peracaula (2018); Benisty (2021); Wright et al. (2020); Nunes and Vagnozzi (2021); Nguyen et al. (2023); Adil et al. (2023). Interestingly, upcoming RSD data from DESI will cover a redshift range up to $z\sim 2.1$, extending by $\sim 0.6$ units the range covered by past surveys, and with small relative uncertainties of $3-6\%$ in the intermediate range $z\in(0.4,1.6)$ Adame et al. (2024b). These new data will allow us to better assess the status of the growth tension.
@@
In this work, we aim to compress the information contained in RSD data into variables that parameterize deviations from the standard growth of perturbations in the late universe, with the aid of a couple of sensible priors. These variables should have a crystal-clear physical interpretation and be applicable to a broad range of modified gravity and dark energy models. More concretely, in this paper, we model the gravitational coupling as a step function and introduce 2 or 3 parameters (depending on the case under study) representing the step widths of the functions. We demonstrate that this model can reproduce the growth of density fluctuations of effective field theories of DE (EFTofDE) with observational accuracy. Other models, such as Galileon gravity, Dvali-Gabadadze-Porrati (DGP) gravity, and $f(R)$ gravity, have already been examined in Denissenya and Linder (2017). We apply the compression technique proposed in Denissenya and Linder (2017) and study its robustness and universality by applying it to three different models within the large class of EFTofDE. In most works in the literature, the time dependence of the EFT parameters or the effective gravitational coupling is assumed when deriving or forecasting constraints from observations Zhao et al. (2009); Ade et al. (2016); Ferté et al. (2019); Abbott et al. (2019); Sakr and Martinelli (2022); Abbott et al. (2023); Casas et al. (2023). These constraints strongly depend on the assumed functional form. In Casas et al. (2017), Casas et al. introduced 6 bins at $30<z<$ 3 and performed forecasts for the constraints from CMB, galaxy clustering and weak lensing. Our focus in this work is galaxy clustering and we show that 2-3 bins are already enough to describe a wide variety of models. More general (continuous) reconstructions of the gravitational coupling can be performed using more sophisticated techniques, such as a Bayesian reconstruction or Gaussian Processes Raveri et al. (2023); Pogosian et al. (2022); Ruiz-Zapatero et al. (2022); Mu et al. (2023), or even Principal Component Analysis (PCA) Pogosian et al. (2010); Hojjati et al. (2012); Asaba et al. (2013), but the resulting constraints cannot be easily projected onto constraints on the parameters of specific models. The PCA approach uses a large number of bins to identify and extract the data’s well-constrained modes, which are insensitive to the number of bins used in the analysis. In principle, a theory prediction can be mapped to these principal component modes. However, because the latter depends on data covariance, this projection is data-dependent. Of course there are models that will not be well described by only 2-3 bins. In this case, the use of more bins or alternative methods as PCA are complementary to our approach. See also Zheng et al. (2024) for a very model-independent method to constrain deviations from standard growth that does not make use of RSD data.
@@
The 2-3 parameter description proposed in Denissenya and Linder (2017) offers a quite model-independent way to reveal the physical characteristics of the underlying theory required to explain the RSD data. It relies on the cosmological principle and the covariant conservation of matter. It can account, e.g., for phenomena such as the suppression of growth at late times Macaulay et al. (2013); Joudaki et al. (2018); Gómez-Valent and Solà (2017); Nesseris et al. (2017); Gómez-Valent and Solà Peracaula (2018); Benisty (2021); Wright et al. (2020); Nunes and Vagnozzi (2021); Nguyen et al. (2023); Adil et al. (2023). Additionally, we will see that it provides an efficient method for constraining EFT parameters with specific time dependencies by projecting the compressed information onto the latter.
@@ IV.2 Current LSS data
In Table 1, we present the list with the 15 RSD data points employed in Sec. VI.1, together with the corresponding references. They are provided by different galaxy surveys and cover the redshift range $z\in(0.01,1.5)$. This sample is quite conservative, as it avoids double-counting issues. For example, CMASS and WiggleZ use different tracers: CMASS focuses on luminous, primarily red galaxies, whereas WiggleZ targets emission-line galaxies in low-to-intermediate mass halos. Although the two surveys overlap in both sky area and redshift, they remain free from double-counting problems Marín et al. (2016). CMASS, in fact, covers a much larger area - $10^{5}$ deg2 compared to WiggleZ’s 816 deg2. Similar considerations apply to the other data points listed in Table 1.
@@ V.1 Methodology
    The latter is obtained in $\Lambda$CDM using the DESI BAO data, the CMB temperature and polarization data from Planck and the CMB lensing data from Planck and ACT Adame et al. (2024a). The results of the fitting analyses carried out using Eq. (15) are displayed in Tables 3-6. It is important to remark that data on $f\sigma_{8}$ alone cannot break the degeneracy between $\Omega_{m}$ and $\mu$, which is apparent from Eq. (2). This is why we impose a prior on the matter density parameter obtained from background observations. In this work, we are interested in the effect of varying $\Omega_{m}$ within the limits of its current background constraints on the constraints on $\mu$. If we avoided the use of a prior on $\Omega_{m}$, we would be only capable of constraining the product $\Omega_{m}\mu$ Zheng et al. (2024).
@@
    with $A_{i},B_{i},D_{i}$ fitting parameters, is accurate enough, since it allows us to recover the constraints on EFT parameters at the posterior level. We illustrate this in Fig. 1, where it is apparent that the expected uncertainties of the final DESI data are much larger than the error of the fitting formula. See also Sec. V.2.
@@
    where $\boldsymbol{p}(c_{j},\Omega_{m})$ and $\boldsymbol{p}_{\mathrm{mean}}$ are the vectors containing the various $\mu_{i}$’s and $\Omega_{m}$ from theory and data, respectively. The covariance matrix $Cov$ is also obtained in point 2. With a prior on $\Omega_{m}$, we found that the Gaussian approximation for these parameters is valid for the DESI mock data. We have explicitly tested the validity of the Gaussian approximation by using the full posterior of the $\mu_{i}$’s and comparing the results with those obtained using the Gaussian approximation. We found negligible differences compared to the typical uncertainties. Thus the use of the Gaussian approximation is fully justified, since it simplifies the analysis and does not induce any significant error.
@@ V.2 Results
Figure 1: Comparison of the theoretical curves of $f(z)\sigma_{8}(z)$ obtained for the various EFT models employed in this paper with $c_{B}=0.4$ and $\Omega_{m}=0.3$ (in blue), and the corresponding best-fit curves of the bin model (in red or orange). The latter have been obtained following the procedure explained in point 3 of Sec. V.1. We also show the DESI mock data, as described in Sec. IV.1. The differences between the theoretical and fitted curves are negligible compared to the forecasted uncertainties. In the right column we plot the percentage relative difference $\Delta f\sigma_{8}(z)/f\sigma_{8}(z)$ [%] between the theoretical and best-fit results. They are smaller than $\sim 1.5\%$ at $z<0.5$ and smaller than $0.2\%$ at $z>0.5$ in absolute value, with the only exception of the fit of the propto-Scale model using $\mu$-2param. As explained in Sec. V.2.3, in this case we need 3 parameters to improve the accuracy of the fitting formula. The sudden change in the derivative of $f(z)\sigma_{8}(z)$ found at $z=1$, which is apparent in the plots on the right, is expected and caused by the discontinuity of $\mu(z)$ at this redshift, cf. formulas (3) and (4)
@@
This is the covariance matrix that enters Eq. (20) and is employed in the second Monte Carlo run to obtain the projected constraints on the EFT parameters in the bin method. Figure 2 demonstrates that the Gaussian approximation is sufficient, since the two-dimensional contours are very close to perfect ellipses. While we present the mean values and errors of the $\mu_{i}$’s from all analyses in this work in the various tables, we provide the covariance matrix only for this particular case to maintain compactness.
@@ V.2.2 Inv-Hubble-Squared
[…]
@@
Table 6: Same as Table 3, but for the two propto-Scale models studied in this paper. On the top we show the results obtained considering $c_{B}\neq 0$ and $c_{M}=0$, while on the bottom we display the results obtained when $c_{B}=-c_{M}$. We use a $\Lambda$CDM background. In the last row, we quantify the discrepancy between the values of $c_{B}$ obtained from the direct and binning methods making use of Eq. (22). In all cases $T<0.17$, which demonstrates the good performance of our approach.
@@ VI.1 Current data
$\mu_{\rm 2p}(a)=1+\frac{1}{2}\left[\mu_{2}+\left(\frac{\mu_{1}-\mu_{2}}{2}\right)\left(1+\tanh\{\beta(2-1/a)\}\right)-1\right]\times\left[1+\tanh\{\beta(4-1/a)\}\right]\,.$ (23)
@@
Figure 3: Example of $\mu_{\rm 2p}(z)$ with $\mu_{1}=1.3$ and $\mu_{2}=1.1$, and of $\mu_{\rm 3p}(z)$ with $\mu_{1}=1.1$, $\mu_{2}=0.9$ and $\mu_{3}=1.2$. In both cases we have set $\beta=5$. See Eqs. (23) and (24).
$\mu_{\rm 3p}(a)=1+\frac{1}{2}(\tilde{\mu}(a)-1)\left[1+\tanh\{\beta(10-1/a)\}\right]\,,$ (24)
@@
A couple of examples of $\mu_{\rm 2p}(z)$ and $\mu_{\rm 3p}(z)$ are provided in Fig. 3.
@@
Our results are provided in Table 7 and Fig. 4. In Fig. 5 we compare the best-fit curves of $f(z)\sigma_{8}(z)$ obtained with $\mu$-2param, $\mu$-3param and the $\Lambda$CDM. We find $\mu_{1}<1$ at $\sim 2.6-2.8\sigma$ C.L. with both parametrizations. The data on $f(z_{i})\sigma_{8}(z_{i})$ prefer a lower amount of structure formation in the universe than in $\Lambda$CDM at $z<1$. This explains why our parametrizations are capable of alleviating the tension and decreasing significantly the values of $\chi^{2}_{\rm min}$. Notice that 13 out of the 15 data points on $f(z_{i})\sigma_{8}(z_{i})$ employed in the analysis are inside that redshift range. These results are aligned with those reported in previous works in the literature, see e.g. Macaulay et al. (2013); Joudaki et al. (2018); Gómez-Valent and Solà (2017); Nesseris et al. (2017); Gómez-Valent and Solà Peracaula (2018); Benisty (2021); Wright et al. (2020); Nunes and Vagnozzi (2021); Nguyen et al. (2023); Adil et al. (2023) and references therein. At redshifts $z>1$ we only have two data points, with larger central values than in $\Lambda$CDM. This leads to values of $\mu_{2}$ in the parametrization $\mu$-2param slightly greater than $1$, $\mu_{2}=1.10\pm 0.09$, but compatible with the standard value ($\mu=1$) at $\sim 1\sigma$ C.L. In $\mu$-3param we find a significant degeneracy in the plane $\mu_{2}-\mu_{3}$ (cf. Fig. 4) due to the lack of data points at $z>1$. Current data are not able to set strong constraints on these two parameters simultaneously, and actually the posterior distributions hit the prior boundaries.
@@
Figure 5: Best-fit curves of $f(z)\sigma_{8}(z)$ obtained using the two parametrizations $\mu_{\rm 2p}(z)$ and $\mu_{\rm 3p}(z)$, Eqs. (23) and (24), together with the data points of Table 1 and the curve obtained using the $\Lambda$CDM best-fit model.
@@ VII Conclusions
Although we focused on obtaining the constraints on $\mu$ using the compressed observable $f\sigma_{8}$ in this paper, our method can also be eventually applied to data on $f\sigma_{12}$ Sánchez (2020); Semenaite et al. (2022, 2023), in case they become available in the future. The 2-3 parameter description of $\mu$ can be also implemented in the EFTofLSS approach to obtain the constraints on $\mu_{i}$ directly. It is known that the EFTofLSS approach suffers from prior projection effects due to a large number of nuisance parameters. Ref. Piga et al. (2023) tested a specific model, which gives a scale independent linear modification as in EFTofDE models considered in this paper and found a strong projection effect, i.e. even if the synthetic data is created by $\Lambda$CDM, the 1D marginalised constraint strongly favours non-zero modified gravity parameter after marginalising over nuisance parameters of EFTofLSS. We expect to see a similar effect for $\mu$ and this would become more severe with the increased number of parameters as they lead to more degeneracies among $\mu$ in different bins and the primordial amplitude. A fewer bins would alleviate this problem. The 2-3 parameter description is ideal as we do not need to assume a specific time dependence of $\mu$, and it is general enough to cover a wide variety of models as shown in this paper and Denissenya and Linder (2017). It remains to see whether the same binning of $\mu$ is enough for other LSS observables such as CMB (ISW, lensing) and weak lensing. This will be addressed in future work.